OpenStudy (prepfortears):

Will award medal Find an exact value cos((19pi)/12)

OpenStudy (anonymous):

Look at your unit circle, did you notice how this value is not a part of what is labeled?

OpenStudy (anonymous):



OpenStudy (anonymous):

Every time you go around the unit circle it is 2pi.

OpenStudy (anonymous):

Your value is greater than 2pi. \[\frac{ 19\pi }{ 2 }~=~9\pi~and~\frac{ \pi }{ 2 }\](as a mixed fraction)

OpenStudy (anonymous):